# Ralph Abbey1, Jorge Silva1, Emiliano De Cristofaro3 1SAS   2UCL   3UC Riverside

### v2

@@ Abstract
The Synthetic Minority Over-sampling Technique (SMOTE) is one of the most widely used methods for addressing class imbalance and generating synthetic data. Despite its popularity, little attention has been paid to its privacy implications; yet, it is used in the wild in many privacy-sensitive applications. In this work, we conduct the first systematic study of privacy leakage in SMOTE: We begin by showing that prevailing evaluation practices, i.e., naive distinguishing and distance-to-closest-record metrics, completely fail to detect any leakage and that membership inference attacks (MIAs) can be instantiated with high accuracy. Then, by exploiting SMOTE’s geometric properties, we build two novel attacks with very limited assumptions: DistinSMOTE, which perfectly distinguishes real from synthetic records in augmented datasets, and ReconSMOTE, which reconstructs real minority records from synthetic datasets with perfect precision and recall approaching one under realistic imbalance ratios. We also provide theoretical guarantees for both attacks. Experiments on eight standard imbalanced datasets confirm the practicality and effectiveness of these attacks. Overall, our work reveals that SMOTE is inherently non-private and disproportionately exposes minority records, highlighting the need to reconsider its use in privacy-sensitive applications.
@@ 1 Introduction
Synthetic Data. SMOTE has also gained traction as a method for generating synthetic tabular data. Often used as a benchmark for more advanced models like GANs and VAE, SMOTE has been shown to perform on par with, or even better than, generative approaches [39, 32]. Moreover, it is being used as a reference point for diffusion-based models published at top-tier machine learning conferences [34, 59, 47, 42]. SMOTE remains one of the most widely applied algorithms for medical synthetic data [30], and beyond machine learning research, it has also been recognized as a promising technique for improving access to census data by public sector entities [45, 1].
@@
Our experiments on eight standard imbalanced datasets, summarized in Table 1, demonstrate that:
@@ 3 Related Work
SMOTE has recently served as a baseline in several diffusion-based generative models [34, 59, 47, 42], all published at top-tier machine learning venues. These studies primarily rely on the Distance to Closest Record (DCR) [61] as a privacy proxy, consistently showing smaller distances for SMOTE and claiming this as evidence of weaker privacy. Kotelnikov et al. [34] additionally employ the “full black-box” attack from [8], which still reduces to DCR as its core signal.
More recently, this view has been challenged by Sidorenko and Tiwald [54], who show that some diffusion models [34, 42] achieve lower DCR values than SMOTE, thereby leaking more information about the training data. However, DCR itself has been shown to be an inadequate privacy metric, which consistently underestimates privacy leakage [28, 4, 19, 57]. To the best of our knowledge, SMOTE has not yet been systematically evaluated with state-of-the-art MIAs or through reconstruction attacks, leaving a critical gap in understanding its privacy risks.
@@ 4.1 Adversarial Model
In other words, all features are continuous and no three points lie on the same line. These assumptions are reasonable for (high-dimensional) continuous data and hold for all datasets in our experiments (see Section 5). Moreover, assuming $k\geq 3$ is necessary to uniquely identify the intersection point of the lines formed by vectors connecting neighboring points, and is a reasonable choice in practice.
@@ 17    Add $N(x_{m})\cap C^{1}$ to queuequeue
Complexity. The nearest-neighbor search is the main cost. With brute-force search ($\mathcal{O}(nd)$ per query), each of the $n$ records requires finding $kr$ neighbors ($\mathcal{O}(nd)$) and checking all neighbor pairs ($\mathcal{O}((kr)^{2})$), yielding a worst-case complexity of $\mathcal{O}(n^{2}d+n(kr)^{2})$. In practice, optimized methods (e.g., KD/ball trees) reduce search to $\mathcal{O}(\log n)$ for small $d$ ($d\leq 32$ in our datasets). Also, since $k$ and $r$ are typically (small) constants, the effective complexity is much lower, and the algorithm completes in under three minutes on all tested datasets.
Accuracy Analysis. We now analyze the expected performance of DistinSMOTE.
@@
Under Assumptions 12, the labeling rule proposed by the DistinSMOTE attack achieves perfect precision and recall.
@@
DistinSMOTE follows this labeling rule: for such lines, it marks interior points as synthetic and endpoints as real, while any real point not on such a line (i.e., not used in interpolation) is also labeled real. This ensures all synthetic points are eliminated and all real points are preserved. This leads to perfect precision and recall (“real” is the positive class). ∎
@@ 4.3 Reconstruction Attack
Complexity. The worst-case time complexity is very similar to DistinSMOTE, i.e., $\mathcal{O}(n^{2}d+n(kr)^{2})$. While there are two additional factors, namely, $\mathcal{O}(nkr)$ for checking neighbors after identifying a collinear triplet and $\mathcal{O}(n^{2})$ for finding line intersections, these are dominated by existing terms and can be ignored. The practical complexity is much lower, and the attack runs in at most three minutes on all datasets.
Accuracy. Next, we analyze ReconSMOTE’s expected performance.
Theorem 2 (ReconSMOTE perfect precision).
Under Assumptions 13, the records reconstructed by ReconSMOTE are guaranteed to be real, i.e., the attack achieves perfect precision.
Proof. By SMOTE construction, each synthetic point lies on a segment strictly between two real endpoints. Thus, every detected line in $X^{1}_{syn}$ corresponds uniquely to a pair of real records.
Under the global non-collinearity assumption, the only point lying on three or more such lines is the shared real endpoint. Since ReconSMOTE retains only intersections supported by at least three distinct lines (as $k\geq 3$), every reconstructed point must be a true real point. Hence, precision is $1$. $\square$
Theorem 3 (ReconSMOTE expected recall (approximate)).
$\mathbb{E}[\mathrm{Recall}]\geq$
$\max\!\left\{0,\ \frac{k\!\left(1-e^{-\lambda}\!\left(1+\lambda+\tfrac{\lambda^{2}}{2}\right)\right)-2}{k-2}\right\}.$
Proof. At each generation step, SMOTE first selects a minority record $x_{i}$ uniformly from the $n_{1}$ available, and then one of its $k$ nearest neighbors $x_{j}$ uniformly. Thus, each of the $n_{1}k$ possible minority-neighbor (directed) segments is chosen with probability $\frac{1}{n_{1}k}$ at every step.
Let $C_{ij}$ denote the number of synthetic points generated on segment $(x_{i},x_{j})$. Since each of the $n_{0}-n_{1}$ synthetic points is assigned independently to a segment with probability $\frac{1}{n_{1}k}$, we have $C_{ij}\sim\mathrm{Binom}(n_{0}-n_{1},\frac{1}{n_{1}k})$. For analytic tractability, we approximate this by $\mathrm{Poisson}(\lambda)$ with mean $\lambda=\frac{n_{0}-n_{1}}{n_{1}k}$. This is standard when $n_{0}-n_{1}$ is large and $\frac{1}{n_{1}k}$ is small, which holds in practice.
@@ 30  Add $x^{*}$ to $R^{1}$ and record support line
This is the probability that $x_{i}$ is identifiable. Since recall is the fraction of minority records that are identifiable, its expectation equals the average of these probabilities over all $n_{1}$ records. Because each $x_{i}$ is treated symmetrically in SMOTE and we look at directed segments, the average equals the bound derived above. Hence we obtain the stated lower bound on the expected recall. $\square$
@@
For brevity and clarity, Theorem 3 is derived under the Poisson approximation with $C_{ij}$ counted only in one direction. We refer to this as the approximate bound. In Appendix A, we provide a more detailed analysis and derive a tighter bound without these simplifications; we call it the exact bound. Finally, in Appendix B, we visualize the differences between the bounds under various conditions.
@@ 5 Experimental Evaluation
Figure 2: Datasets overview, where $r$ denotes the imbalance ratio ($n_{0}/n_{1}$), $n$ the number of records, and $d$ the number of features.
Datasets. We conduct experiments on eight standard imbalanced datasets, each with a binary classification task, obtained from the imblearn library [37] (originally from the UCI ML Repository) and used in prior work [12, 52] These datasets vary quite significantly in size (336 to 11,183 records), dimensionality (6 to 32 features), imbalance ratios (8.6 to 130), and prediction task (target), as shown in Table 2.
Implementations. We use the imblearn [37] implementation of SMOTE and sklearn [48] for machine learning classifiers. Both DistinSMOTE and ReconSMOTE are highly efficient, running in under three minutes on any dataset on an Apple M4 MacBook with 24GB RAM. The naive methods are similarly fast, while the MIAs require up to 30 minutes per dataset. We will release the source code for our attacks along the final version of the paper.
@@ 5.1 Augmented Data
Membership Inference. Next, we evaluate MIAs [53, 6] using the repeated classification game from Section 2. For a given target record, we train 200 classifiers (a multi-layer perceptron with two hidden layers) on augmented datasets generated via SMOTE: half of the training datasets include the target record, and half exclude it. We then use the classifiers’ predictions on the target to simulate an adversary’s confidence in distinguishing membership, and calculate AUC. Following prior work [58, 24], we train target-specific attacks in a leave-one-out setting, which provides a more accurate estimate of privacy leakage. This procedure is repeated for 100 randomly selected targets (or all minority records), and we report the average. Overall, this requires training roughly 20k SMOTE models and classifiers per dataset.
@@ 6 Conclusion
Our work highlights the fundamental privacy limitations of SMOTE [7], one of the most widely adopted techniques for improved learning on imbalanced data. The effectiveness of our novel, near assumption-free attacks (DistinSMOTE and ReconSMOTE), demonstrates that real minority records – precisely the ones SMOTE aims to better represent – are exposed to significant privacy risk. Nonetheless, SMOTE remains an effective and easy-to-use technique in non-privacy-sensitive applications where utility is the primary concern. We are confident our findings will be valuable to researchers and practitioners deploying solutions that process or release sensitive data, and will motivate them to adopt more robust privacy-preserving techniques.
Limitations and Future Work. Our attacks currently operate on continuous data and are tested only on the original SMOTE implementation. However, our findings generalize to many SMOTE variants – such as BorderlineSMOTE [25], ADASYN [27], SVMSMOTE [44], and cluster/hybrid-based methods, etc. – as they all rely on line-segment interpolation to generate synthetic samples. Extending our attacks and defenses to these variants is a promising direction for future work.
@@
Ethics Statement. Our work does not involve attacking live systems or private datasets. Our goal is to demonstrate the importance of emphasizing privacy considerations and relying on established notions of privacy when processing sensitive, imbalanced data in critical domains.

### v3

@@ Abstract
The Synthetic Minority Over-sampling Technique (SMOTE) is one of the most widely used methods for addressing class imbalance and generating synthetic data. Despite its popularity, little attention has been paid to its privacy implications; yet, it is used in the wild in many privacy-sensitive applications. In this work, we conduct the first systematic study of privacy leakage in SMOTE: we begin by showing that prevailing evaluation practices, i.e., naive distinguishing and distance-to-closest-record metrics, completely fail to detect any leakage and that membership inference attacks (MIAs) can be instantiated with high accuracy. Then, by exploiting SMOTE’s geometric properties, we build two novel attacks with very limited assumptions: DistinSMOTE, which perfectly distinguishes real from synthetic records in augmented datasets, and ReconSMOTE, which reconstructs real minority records from synthetic datasets with perfect precision and recall approaching one under realistic imbalance ratios. We also provide theoretical guarantees for both attacks. Experiments on eight standard imbalanced datasets confirm the practicality and effectiveness of these attacks. Overall, our work reveals that SMOTE is inherently non-private and disproportionately exposes minority records, highlighting the need to reconsider its use in privacy-sensitive applications and as a baseline for assessing the privacy of modern generative models.
@@ 1 Introduction
Synthetic Data. SMOTE has also gained traction as a method for generating synthetic tabular data. Often used as a baseline for more advanced models like GANs and VAE, SMOTE has been shown to perform on par with, or even better than, generative approaches [42, 35]. Moreover, its extensive use as a baseline has led to modern diffusion-based models (not explicitly designed with privacy in mind) to be characterized as privacy-preserving simply because they outperform SMOTE, a pattern repeatedly observed in top-tier machine learning publications [37, 63, 50, 45]. SMOTE is also applied for medical synthetic data [33], and beyond machine learning research, it has been recognized as a promising technique for improving access to census data by public sector entities [48, 1].
@@
Our experiments, summarized in Table 1, on eight standard imbalanced datasets demonstrate that:
@@
    SMOTE and DCR are unreliable: evaluating SMOTE with privacy metrics like DCR gives a misleading assessment of its privacy and should not be used to validate other generative models.
  4. 4.
@@ 3 Related Work
SMOTE has recently served as a baseline in several diffusion-based generative models [37, 63, 50, 45], all published at top-tier machine learning venues. A common pattern across these studies is the “SMOTE + DCR” workflow: they rely on the Distance to Closest Record (DCR) [65] as the primary privacy proxy, consistently reporting smaller DCR values for SMOTE and interpreting this as evidence that the newly proposed models are privacy-preserving. Kotelnikov et al. [37] additionally employ the “full black-box” attack from [8], which still reduces to DCR as its core signal.
More recently, this view has been challenged by Sidorenko and Tiwald [57], who show that some diffusion models [37, 45] actually achieve lower DCR values than SMOTE, thereby leaking more information about the training data. However, DCR itself has been shown to be an inadequate privacy metric – it consistently underestimates leakage [31, 4, 22] and does not correlate with leakage detected by MIAs [61]. To the best of our knowledge, despite its prominent role as a baseline, SMOTE has not yet been systematically evaluated with state-of-the-art MIAs or any model-specific attacks. This leaves a critical gap in understanding SMOTE’s true privacy risks and calls into question the validity of privacy claims across a recent line of generative-model research.
@@ 4.1 Adversarial Model
In other words, all features are continuous and no three points lie on the same line. These assumptions align naturally with (high-dimensional) continuous data, are non-restrictive in practice, and, crucially, are satisfied by all datasets in our main experiments (see Section 5). Moreover, assuming $k\geq 3$ is necessary to uniquely identify the intersection point of the lines formed by vectors connecting neighboring points, and is a standard, practical choice in typical SMOTE configurations (default is 5).
@@ 17    Add $N(x_{m})\cap C^{1}$ to queuequeue
Complexity. The nearest-neighbor search is the main cost. With brute-force search ($\mathcal{O}(nd)$ per query), each of the $n$ records requires finding $kr$ neighbors ($\mathcal{O}(nd)$) and checking all neighbor pairs ($\mathcal{O}((kr)^{2})$), yielding a worst-case complexity of $\mathcal{O}(n^{2}d+n(kr)^{2})$. In practice, optimized methods (e.g., KD/ball trees) reduce search to $\mathcal{O}(\log n)$ for small $d$ ($d\leq 32$ in our main datasets). Also, since $k$ and $r$ are typically (small) constants, the effective complexity is much lower, and the algorithm completes in under three minutes on all main datasets.
Accuracy Analysis. We analyze DistinSMOTE, with the theorem below formalizing the theoretical correctness of its labeling rule, achieving perfect precision and recall.
@@
Under Assumptions 12, the labeling rule in the DistinSMOTE attack achieves perfect precision and recall (with probability 1).
@@
DistinSMOTE follows this labeling rule: it finds lines via local search (lines 9-11 of Algorithm 2) and marks interior points as synthetic and endpoints as real (lines 12-13), while any real point not on such a line (i.e., not used in interpolation) is also labeled real. The local search guarantees all points are visited efficiently, and the labeling rule ensures all synthetic points are removed and all real points preserved. This leads to perfect precision and recall (with probability 1, except for negligible numerical precision effects that do not occur in practice). ∎
@@ 30  Add $x^{*}$ to $R^{1}$ and record support line
Complexity. The worst-case time complexity is very similar to DistinSMOTE, i.e., $\mathcal{O}(n^{2}d+n(kr)^{2})$. While there are two additional factors, namely, $\mathcal{O}(nkr)$ for checking neighbors after identifying a collinear triplet and $\mathcal{O}(n^{2})$ for finding line intersections, these are dominated by existing terms and can be ignored. The practical complexity is much lower, and the attack runs in at most three minutes on all main datasets.
Accuracy Analysis. Next, we analyze ReconSMOTE; the following theorems give theoretical guarantees for its reconstruction rule, achieving perfect precision and a lower bound on expected recall.
Theorem 2 (ReconSMOTE perfect precision).
Under Assumptions 13, the records reconstructed by ReconSMOTE are guaranteed to be real, i.e., the attack achieves perfect precision (with probability 1).
Proof.
By SMOTE construction, each synthetic point lies on a segment strictly between two real endpoints. Accordingly, every detected line in $X^{1}_{syn}$ (obtained from collinear sets identified by the local search in lines 8-17 of Algorithm 3) corresponds uniquely to a pair of real records. The local search correctly groups all synthetic points from that pair into a single collinear set without introducing spurious collinearities.
Under the global non-collinearity assumption, the only point lying on three or more such lines is their shared real endpoint. ReconSMOTE exploits this by intersecting detected lines (lines 23-25) and retaining only points supported by at least three distinct lines (line 27). Since synthetic points lie on exactly one SMOTE line, whereas real endpoints lie on three or more, any retained intersection must be a true real point. Hence, every reconstructed record is real, so precision is $1$ (with probability 1). ∎
For clarity and tractability, the next theorem uses a simplified formulation with directed edges and a Poisson approximation, giving a conservative bound (the approximate bound) that ignores overlapping neighbor relations in the SMOTE graph.
Theorem 3 (ReconSMOTE expected recall (approximate)).
$\mathbb{E}[\mathrm{Recall}]\geq\max\!\left\{0,\ \frac{k\!\left(1-e^{-\lambda}\!\left(1+\lambda+\tfrac{\lambda^{2}}{2}\right)\right)-2}{k-2}\right\}.$
Proof.
At each generation step, SMOTE first selects a minority record $x_{i}$ uniformly from the $n_{1}$ available, and then one of its $k$ nearest neighbors $x_{j}$ uniformly. Thus, each of the $n_{1}k$ possible minority-neighbor (directed) segments is chosen with probability $\frac{1}{n_{1}k}$ at every step.
Let $C_{ij}$ denote the number of synthetic points generated on segment $(x_{i},x_{j})$. Since each of the $n_{0}{-}n_{1}$ synthetic points is assigned independently to a segment with probability $\frac{1}{n_{1}k}$, the vector of all $C_{ij}$ follows a multinomial distribution with $\sum_{ij}C_{ij}=n_{0}{-}n_{1}$, and each $C_{ij}$ is marginally distributed as $\mathrm{Binom}(n_{0}{-}n_{1},\frac{1}{n_{1}k})$. For analytic tractability, we approximate this by $\mathrm{Poisson}(\lambda)$ with mean $\lambda=\frac{n_{0}-n_{1}}{n_{1}k}$. This is standard when $n_{0}{-}n_{1}$ is large and $\frac{1}{n_{1}k}$ is small, which holds in practice.
@@
This is the probability that $x_{i}$ is identifiable. Since recall is the fraction of minority records that are identifiable, its expectation equals the average of these probabilities over all $n_{1}$ records. Because each $x_{i}$ is treated symmetrically in SMOTE and we look at directed segments, the average equals the bound derived above. Hence we obtain the stated lower bound on the expected recall. ∎
@@
In Appendix A, we provide a more detailed analysis and derive a tighter bound without the simplifications of Theorem 3; we call it the exact bound. Finally, in Appendix B, we visualize the differences between the bounds under various conditions.
@@ 5 Experimental Evaluation
Figure 2: Main datasets overview, where $r$ denotes the imbalance ratio ($n_{0}/n_{1}$), $n$ the number of records, and $d$ the number of features.
Datasets. We conduct our main experiments on eight standard imbalanced datasets, each with a binary classification task, obtained from the imblearn library [40] (originally from the UCI ML Repository) and used in prior work [13, 55] These datasets vary significantly in size (336 to 11,183 records), dimensionality (6 to 32 features), imbalance ratios (8.6 to 130), and prediction task (target), as shown in Table 2.
Implementations. We use the standard imblearn [40] implementation of SMOTE and sklearn [51] for classifiers. Both DistinSMOTE and ReconSMOTE are highly efficient, running in under three minutes on any dataset from Table 2 on an Apple M4 MacBook with 24GB RAM. The naive methods are similarly fast, while the MIAs take up to 30 minutes per dataset. We will release the source code for our attacks along with the final version of the paper.
@@ 5.1 Augmented Data
MIA. Next, we evaluate MIAs [56, 6] using the repeated classification game from Section 2. For a given target record, we train 200 classifiers (a multi-layer perceptron with two hidden layers) on augmented datasets generated via SMOTE: half of the training datasets include the target record, and half exclude it. We then use the classifiers’ predictions on the target to simulate an adversary’s confidence in distinguishing membership, and calculate AUC. Following prior work [62, 27], we train target-specific attacks in a leave-one-out setting, which provides a more accurate estimate of privacy leakage. This procedure is repeated for 100 randomly selected targets (or all minority records), and we report the average. Overall, this requires training roughly 20k SMOTE models and classifiers per dataset.
@@ 5.3 Take-Aways
6 DistinSMOTE and ReconSMOTE with Relaxed Assumptions
In this section, we test the robustness of our attacks, DistinSMOTE and ReconSMOTE, on augmented/synthetic data while relaxing our assumptions one at a time, e.g., using high-dimensional data, mixed-type data, and $k=2$; results are shown in Table 45, and 6 in Appendix C. Additionally, in Appendix C, we evaluate our attacks on perturbed SMOTE datasets (i.e., linear interpolation with added random noise) and provide a heuristic for running the attacks without knowledge of $k$ and $r$.
High-Dimensional Data. First, we test how the attacks scale to two high-dimensional datasets – up to 96,690 records and 50 features (see Table 4 and Appendix C for results and details about the higgs and miniboone datasets, both with $r=25$). Both DistinSMOTE and ReconSMOTE achieve perfect precision and recall and, as analyzed in Section 4, scale well with $n$ and $d$, taking no more than 8 minutes per dataset, which is highly practical.
Mixed-Type Data. Next, we show that our attacks also apply to mixed-type data, relaxing Assumption 1; see Table 5 and Appendix C for details on the cardio and churn datasets, each containing an equal mix of numerical/categorical features, and with $r=25$. To generate data, we use SMOTE-NC [7] – introduced in the original SMOTE paper and designed for mixed data – via the standard imblearn implementation [40]. For the attacks, we simply ignore categorical features and operate on the continuous ones. As before, we obtain perfect precision and very high recall on both datasets. This approach can also be applied to one-hot encoded data.
SMOTE with $\mathbf{k=2.}$ Finally, we evaluate the attacks on the eight main datasets using SMOTE with $k=2$, relaxing Assumption 3 (see Table 6 in Appendix C). The performance of DistinSMOTE remains unaffected, achieving perfect precision and recall, as expected. As for ReconSMOTE, average recall drops to 0.52 (a 39% decrease compared to SMOTE with $k=5$ in Table 3) because each real record participates in fewer lines, making it harder to reach the support threshold; only records that serve as neighbors to other records beyond their two closest neighbors can be successfully reconstructed. Precision remains perfect as all reconstructed points are still accurate. Nevertheless, reconstructing half the real minority records with perfect confidence is a serious privacy breach.
@@ 7 Conclusion
Our work highlights the fundamental privacy limitations of SMOTE [7], one of the most widely adopted techniques for improved learning on imbalanced data. The effectiveness of our novel, near assumption-free attacks (DistinSMOTE and ReconSMOTE), demonstrates that real minority records – precisely the ones SMOTE aims to better represent – are exposed to significant, previously underestimated privacy risk. Importantly, this also shows that using SMOTE as a baseline with DCR to evaluate privacy is unreliable and can provide a false sense of security. Nonetheless, SMOTE remains an effective and easy-to-use technique in non-privacy-sensitive applications where utility is the primary concern. We are confident our findings will be valuable to researchers and practitioners deploying solutions that process or release sensitive data, motivate them to avoid SMOTE as a privacy benchmark, and encourage them to adopt more robust privacy-preserving techniques.
Limitations and Future Work. Our attacks currently operate on continuous data and are primarily tested on the original SMOTE implementation. While certain numerical instabilities/edge cases are theoretically possible (e.g., a synthetic point appearing collinear with two unrelated real points), their probability is effectively zero in high-dimensional datasets with high numerical precision, and we did not observe any such case in our experiments. Additionally, our findings generalize to many SMOTE variants – such as BorderlineSMOTE [28], ADASYN [30], SVMSMOTE [47], and cluster/hybrid-based methods [16] – as they all rely on line-segment interpolation to generate synthetic samples. In contrast, our attacks are unlikely to be successful against variants like G-SMOTE [15] and GI-SMOTE [10], which generate synthetic points within regions rather than strictly along lines. Nevertheless, these variants are not inherently privacy-preserving and are still likely to remain vulnerable to MIAs. Extending our attacks and developing robust defenses for such methods is a promising direction for future work.
@@
Ethics Statement. Our work does not involve attacking live systems or private datasets. Our goal is to demonstrate the importance of emphasizing privacy considerations and relying on established notions of privacy when processing sensitive, imbalanced data in critical domains. To make our work reproducible, we release the source code for our attacks and experiments: https://github.com/sascommunities/smote-mirrors.